OpenStudy (anonymous):

A total of 363 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was two times the number of adult tickets sold. How many adult tickets were sold?

OpenStudy (shaik0124):

assume The number of student tickets =x The number of adult tickets =y given info in question is The number of student tickets sold was two times the number of adult tickets sold x=2y x+y=363 since x=2y put this in x+y=363 2y+y=363 3y=363 y=363/3 y=121( which are adult tickets)

OpenStudy (anonymous):

Thanks so much dude !

OpenStudy (shaik0124):

pleasure to help u always